The table below presents the stream of costs and benefits for a resource development project. Use the table below to answer the following questions Year Benefits Costs 0 0 60 2 30 60 100 What is the net present value of the benefits at an interest rate of 5%? a. 118.75 b. 123.8:1 c. 119.217 d. 123.5 e. none of the above What is the net present value of the costs at an interest rate of 5%? a. 120 b. 119.0!5 C. 114.29 d. 117.14 e. none of the above Would you start the project or cancel the project at an interest rate of 5% (assuming you are rational and there are no other costs)? a. yes start the project b. no will not start the project c. cannot determine with the information provided Would you start the project or cancel the project at an interest rate of 10% (assuming you are rational and there are no other costs)? a. yes start the project b. no will not start the project c. cannot determine with the information provided
Answer #1

1.

C

Working note:

NPV of benefits = 30/1.05 + 100/1.05^2 = $119.27

2.

D

Working note:

NPV of costs = 60 + 60/1.05 = $117.14

3.

A

I will start the project as NPV of benefits is greater than than the NPV of costs.

4.

B

Working note:

If R = 10%

Then,

NPV of the porject = 30/1.1 + 100/1.1^2 - (60+60/1.1) = - $4.62

I will not start the project as NPV of the project is negative.